The day of yesterday has been very good until, I decided to switched off the system. High dark count rate occured just then. Today, the background pressure is arround 4.10-e7 mbar (in 3 days), slightly better then yesterday. The polarisation scheme is as followed:
- front MCP : -1000 V -> -14000 V
- rear MCP : +1000 V
- anode : +1350 V
V | 100 | 200 | 300 | 400 | 500 | 600 | 700 | 800 | 900 | 1000 | 1100 | 1200 | 1300 | 1400 | 1500 | 1600 | 1700 | 1800 | 1900 | 2000 | 2100 | 2200 | 2300 |
---|
micro A | 2 | 4 | 7 | 9.5 | 11.5 | 14 | 16 | 18 | 20.5 | 23 | 25.5 | 28 | 30 | 32 | 34.5 | 37 | 39 | 41.5 | 44 | 46.5 | |
With respect to yesterday, I can notice that there is 1-2 micro A more above 500 V. Courant leak ?… A first pulse with a repetition rate of 1Hz is detected with 1600 V and the anode at 1300V. The pulse width is about 10 ns and the amplitude around 15-35 mV and is independant of the gage status (on / off). It repetition rate seems to be less with the votlage on the anode around 1100 V. A second pulse with a width <5 ns and and amplitude of 5-10 mV is present for a voltage of 1900 V. At 2000V the darkcount is around 30 Hz while yesterday it around a few Hz !
The system had a strange comportement during the morning and finally the reference channel (D16) died during lunch ! In the afternoon, something had changed, the pictures where much better and we could start to increase the DS16 threshold. The detector was still noisy around 500Hz (+/-300Hz) , and we let it run the whole night at 2100V.